OpenStudy (anonymous):

prove that there is no time when the three hands are equally spaced around the clock

OpenStudy (anonymous):

ie the hands are equally \[2\pi/3\] away from each other

ganeshie8 (ganeshie8):



OpenStudy (anonymous):

hands,A short thick "hour" hand; A long, thinner "minute" hand; A very thin "sweep" seconds hand

OpenStudy (anonymous):

revolutions made by each hand hour rev:day 1:12 hours minute 1:1hour sweep 1:1 min=1/60hours

OpenStudy (anonymous):

let h ,m,s be the angles produced after time t,where t is in hours

OpenStudy (anonymous):

i think we can visualise this as a race to cover a straight line of \[2\pi m\]of length

OpenStudy (anonymous):

and the speeds of all of them will be 1/12 rev per hour,1 rev per hour,60 rev per hour

OpenStudy (anonymous):

at a certain time t they are equally away from each other

OpenStudy (anonymous):

so lets say hour hand has already went y meters so since the sweep hand is the largest then the distance covered by each is s: y+2x=60 rph*t m: y+x=1 rph*t h: y=1/12 rph*t

OpenStudy (anonymous):

provided they are x meters away from each other

OpenStudy (anonymous):

\[2\pi-(2x+y)=x \implies 2\pi=3x+y\\ y=2\pi-3x\]

OpenStudy (anonymous):

so we have to show that there is no such a tym

OpenStudy (anonymous):

\[3y+3x=(60+1+\frac{1}{12})t\\6\pi-9x+3x=6\pi-6x=(\frac{733}{12} )t\] \[t=72\frac{\pi-x}{733} \]

OpenStudy (anonymous):

so at @Hero wat time is this...where they are all equal

OpenStudy (anonymous):

equally diplaced from each other

ganeshie8 (ganeshie8):

one unknown, with two equations, you cant have a solution all the time.

ganeshie8 (ganeshie8):

getting delayed for dinner, brb... ive used angular velocity equations in the wolfram solution above...

OpenStudy (anonymous):

okay guys let me put the original question

OpenStudy (anonymous):

in the original question it says this should be obvius and easy to prove ...this was not the question at hand,i jus wanted to show this easy implicit suppossion

OpenStudy (anonymous):

i wanted to prove it

OpenStudy (anonymous):

why is @Hero laughin...lol

ganeshie8 (ganeshie8):

proof is obvious to see cuz x = 3 2x = 3 equations will not have a solution

OpenStudy (anonymous):

i was concerned wit this part it is well known,and not difficult to prove, that there is no time when the three hands are equally spaced around the clock, with each separating angle 2π/3.

ganeshie8 (ganeshie8):

did u have a look at wolfram link

OpenStudy (anonymous):

let me check it ...

ganeshie8 (ganeshie8):

t = time in seconds, 1) angle between seconds, minutes = 2pi/3 :- t(2pi/60 - 2pi/(60*60)) =2*pi/3 ---------(1) 2) angle between minutes, hours = 2pi/3 :- t(2pi/(60*60) - 2pi/(60*60*60)) = 2*pi/3 -------(2)

OpenStudy (anonymous):

same equation and 60 times of it

ganeshie8 (ganeshie8):

there is no solution for above system

OpenStudy (anonymous):

okay so thats like the ratio we set up.....okay so we assume there is no solution...for this set,hence no such time exist

OpenStudy (anonymous):

wat

ganeshie8 (ganeshie8):

when we say there is no solution for below :- x = 3 2x = 3 are we assuming anything ?

OpenStudy (anonymous):

3! or 360! ways to make the pattern?

ganeshie8 (ganeshie8):

i need to go, cya

OpenStudy (anonymous):

okay,c u later

OpenStudy (anonymous):

so now we will consider the original question i will ask in a new slide